OpenStudy (anonymous):

how do i find a ? log(a)=log(1/2)/15

OpenStudy (anonymous):

is this log base ten?

OpenStudy (anonymous):

yeah, i think... I have to inverse log on both sides to isolate a.... I just don't really know what inverse log is...

OpenStudy (anonymous):

as stupid as that sounds.

OpenStudy (anonymous):

the inverse log is the exponential if your base is ten, then the inverse is \(10^x\)

OpenStudy (anonymous):

you need a calculator for sure, if \[\log(A)=\frac{\log(.5)}{15}\] then \[\large A=10^{\frac{\log(.5)}{15}}\]

OpenStudy (anonymous):

make sure you click on "log base ten" if it is not already there i get about \(.955\)

OpenStudy (anonymous):

@satellite73, hold on

OpenStudy (anonymous):

@satellite73 Yup! I got it now! I can't do the rest of it! Thank you!

OpenStudy (anonymous):

yyw